48. What volume of a 0.65 M solution of glucose (C6H12O6) contains 12.5 g of glucose?
Answer #1

molarity = (W /MW) (1000 / V inmL)

molarity = 0.65 M

W = 12.5 g

MW = 180 g/mol

V = ?

0.65 = (12.5 / 180) (1000 /V)

1000 / V =9.36

V = 106.84 mL

answer = 106.84 mL